Hello I am being asked a question that I do not understand, It asks to Solve a system of equations to find the two numbers described.
The sum of a number and twice a greater number is 8. The sum of the greater number and twice the lesser number is -6

Answer by checkley71(8403)

x+2y=8 or x=-2y+8
2y+2x=-6 now sustitute (-2y+8) for x & solve for y.
2y+2(-2y+8)=-6
2y-4y-16=-6
-2y=-6+16
-2y=10
y=10/-2
y=-5 answer.
x+2*-5=8
x-10=8
x=8+10
x=18 answer.
proof
18+2*-5=8
18-10=8
8=8